OpenStudy (anonymous):

Which is a solution to the equation 5x + 2y = –1? A. (–2, 5) B. (1, –3) C. (2, –4) D. (–1, 4)

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

K

OpenStudy (anonymous):

Ordered pairs are in the form of (x, y). For A, x=2 and -4=y Insert the numbers where its respective variable is. 5(2) + 2(-4) = -1 Then, solve using Order of Operations. 5 * 2 + 2 * -4 = -1 10 + 2 * -4 = -1 10 + -8 = -1 10 - 8 = -1 2 = -1

OpenStudy (anonymous):

Thanks!

OpenStudy (the_fizicx99):

Plug in those values for x and y, see if it gives you a true statement, 5x + 2y = –1 Using A, (-2,5) 5(-2) + 2(5) = -1 -10 + 10 = -1 -10 + 10 = 0 0 = 1 <-- therefore this isn't a true statement.

OpenStudy (anonymous):

It is B

OpenStudy (anonymous):

equation of the line= 5x+2y=1 the equation is of the form ax+by+c=0 then the slope of the line= -(x coefficient)/(y coefficient) step1:1st we convert given line in to ax+by+c=0 for this we subtract 1 on both sides 5x+2y-1=1-1 5x+2y-1=0 now the equation is of the form ax+by+c=0 slope= -(x coefficient)/(y coefficient) here x coefficient=5,y coefficient=2 slope= -(5)/2

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

Phantom, if it's B, then why didn't you put x=1 and y=-3?

OpenStudy (anonymous):

Why cant i?

OpenStudy (anonymous):

they needed helped so i helped

OpenStudy (the_fizicx99):

Its neither A... and because you're taking credit for something that's not yours..

OpenStudy (the_fizicx99):

That typo, not*

OpenStudy (anonymous):

did i say it was mine? no.

OpenStudy (anonymous):

The way Phantom said it the first time, it should be C.

OpenStudy (anonymous):

im not taking credit because i am it on my homework or anything

OpenStudy (the_fizicx99):

If its C, Then, use (2, –4) 5(2) + 2(-4) = –1 10 + (- 8) = -1 10 - 8 = 2 2 = -1 <-- not true

OpenStudy (anonymous):

Nvm, case closed. XD

OpenStudy (the_fizicx99):

xD lol

OpenStudy (anonymous):

It's B.

OpenStudy (anonymous):

ok then the internet was wrong. so i taking taking credit for something that wasnt right? w.e bye

OpenStudy (anonymous):

Bye

OpenStudy (the_fizicx99):

Yea

OpenStudy (anonymous):

I'm closing the question now.

OpenStudy (anonymous):

ok bye